# What Do You See? Enhancing Zero-Shot Image Classification with Multimodal Large Language Models

### v4

@@ 3 Experiments
In our experiments, we employed Gemini [9] as our multimodal LLM, $g$, for generating image descriptions and initial predictions. We utilized CLIP (ViT-L/14) [21] as our cross-modal embedding encoder models, $f_{i}$ and $f_{t}$, to encode input testing images, image descriptions, and initial predictions generated by Gemini.
@@ 39 Returns:
Appendix B Additional Results
@@
Appendix C Broader Impact
Our work introduces a method for zero-shot image classification that leverages the power of multimodal large language models (LLMs) not only during the classifier model construction phase but also at inference time. We achieve this by generating comprehensive textual representations directly from input images. These representations are then combined with the input images for classification, resulting in a significant enhancement in accuracy.

### v5

@@ 3 Experiments
In our experiments, we employed Gemini [11] as our multimodal LLM, $g$, for generating image descriptions and initial predictions. Results using GPT [3] are provided in the supplemental materials. We utilized CLIP (ViT-L/14) [23] as our cross-modal embedding encoder models, $f_{i}$ and $f_{t}$, to encode input testing images, image descriptions, and initial predictions generated by Gemini.
@@ 39 Returns:
Appendix B Additional Results Using Gemini
@@
Appendix C Results Using GPT
Our primary goal in this paper is not to compare different LLMs, but to evaluate zero-shot classification methods under a fair setting using the same vision and language backbones. Nonetheless, to assess the generalization of our method across different LLMs, we conducted experiments using GPT-4 [3] as the LLM backbone. Results on the ImageNet validation set [5] are reported in Table 8, alongside CuPL [22], using identical backbones (ViT-L/14 for CLIP and GPT for the LLM). We also compare with recent methods [19, 7, 8] that adopt similar experimental setups. As shown, our method continues to achieve state-of-the-art results even with GPT-4. Notably, using only the raw GPT output—used to construct the ‘initial prediction feature’—achieves 61.1% on ImageNet (ROUGE-F1), which underscores the necessity of our approach: LLM-generated outputs do not always align precisely with class labels.
Table 8: Results w/ GPT LLM [3] and CLIP (ViT-L/14) [23] on ImageNet [5]. The best results are highlighted in yellow.
Method Accuracy (%)
CuPL [22] 75.7
CuPL [22] + REAL-Prompt [8] 75.9
DCLIP [19] 74.5
DCLIP + REAL-Prompt [8] 74.9
MPVR (GPT+TEMP) [7] 76.8
[…]
Appendix D Broader Impact
Our work introduces a method for zero-shot image classification that leverages the power of multimodal LLMs not only during the classifier model construction phase but also at inference time. We achieve this by generating comprehensive textual representations directly from input images. These representations are then combined with the input images for classification, resulting in a significant enhancement in accuracy.